# Surface bubble lifetime in the presence of a turbulent air flow, and the effect of surface layer renewal

### v2

@@ 4 Results
Figure 4: (a) Illustrations of thickness variations in the bubble cap using color interferometry. From left to right the Reynolds number is 0, 142, and 466. (b) Lifetime as a function of the integral Reynolds number. Each circle represents a single bubble lifetime, the boxplot adds additional statistical information: (green line) median, (orange dashed line) mean, (box width) interquartile range, (whiskers) furthest data point within 1.5 times the interquartile range from the box. The data corresponding to a fan RPM of 50 or $Re=18$ has been slightly shifted for clarity. The data presented is obtained with 200 $\mu$mol/L of SDS and at a relative humidity of 50 %. (b) Mean lifetime as a function of the parameter for several surfactant concentrations, wind speeds and relative humidities. Color indicates SDS concentration and markers show relative humidity. The dashed line is a power law of slope 0.5.
@@ 4.1 Qualitative observations
The presence of wind strongly disrupts the classical picture of the motion of the patches in the bubble cap (Illustrated with three panels of Fig. 4 (a)). In these frames white light creates interference pattern within the bubble cap and provides a view of thickness variations. In the top panel (no wind) we see patches generated at the foot of the bubble through marginal regeneration and rising in the cap. These patches allow the exchange of fluid between the bubble cap and the bath and are responsible for the drainage of the cap. They are generated through an instability of the pinch at the foot of the bubble (Lhuissier and Villermaux,, 2012; Gros et al.,, 2021; Trégouët and Cantat,, 2021). These patches are thinner than the rest of the film and therefore buoyant, and they detach from the pinch at the foot when they are large enough. Each patch generated therefore corresponds to a loss of volume in the bubble cap, and it has been shown that the volume change due to generation of these patches exactly matches the viscous limited flow out of the cap (Miguet et al.,, 2021). The motion of these patches in the cap can also be taken into account to model flows induced by surface tension gradients either due to evaporation or heating for instance (Poulain et al.,, 2018).
This picture is largely modified by the presence of wind (middle and bottom panel). Instead of rising straight up because of buoyancy, the patches are entrained by the wind in the air close to cap. They are therefore constantly being deformed and pushed around by the strongly intermittent wind close to the bubble. In the middle panel we can still discern the patches, but some of them have a mostly horizontal velocity instead of vertical and their shape is quite different from the classical picture. In the bottom panel (largest wind forcing) the cap is almost completely mixed with the patches now taking the form of horizontal bands in the cap. In the case with the lowest wind forcing ($Re=16$, not shown), the motion of the patches varies between being very similar to the no wind case and being entrained by the wind as an eddy reaches the proximity of the bath surface. We have verified that the presence or absence of the overflow has no significant effect on the bubble cap images or the conclusions above.
@@
Wind strongly reduces the mean lifetime of surface bubbles in solutions of pure surfactants. For a given solution and relative humidity, the lifetime can vary by a factor up to 5 only due to the intensity of the wind forcing. Fig. 4 (b) illustrates this effect, (with 200 µmol/L of SDS and a relative humidity of 50%) where we can see several effects of the presence of wind: each point is an individual bubble lifetime measurement and the box plot shows the most important metrics of the distribution. The mean lifetime (orange horizontal lines) monotonously decreases as the Reynolds number increases, and so does the spread around the mean. The reduction of the spread is reminiscent of our previous study on soap films (Aurégan and Deike,, 2025), where we found that mixing the air avoids unwanted local fluctuations of the humidity field close to the film. A similar process may be happening here as conditions get more homogeneous and statistically invariant in time as the wind forcing is increased. The reduction in the mean lifetime may a priori be due to many effects: early bursting due to a perturbation from the wind, increased drainage through direct evaporation from the cap or increased flow through the bubble foot because of the strong cap mixing for instance. In the following we derive a drainage model taking into account only direct evaporation and show that it is sufficient to explain most of the observed variation of lifetime.
@@ 4.2 Drainage
We model the effect of the wind on bubble lifetime by taking into account its effect on the evaporative rate. The effect of the wind-induced motion in the cap is not considered in the following. There are two mechanisms responsible for the drainage of the cap of the bubbles: viscosity limited flow from the cap towards the bath due to the capillary pressure through the foot of the bubble and direct evaporation from the cap. Following the derivation from previous studies (Lhuissier and Villermaux,, 2012; Poulain et al.,, 2018), the velocity through the foot of the bubble reads $u=\alpha_{d}(\gamma/\mu)(h/R)^{3/2}$, with $\gamma$ and $\mu$ the surface tension and viscosity of the liquid, respectively. $\alpha_{d}$ is non-dimensional a constant of order unity. The evaporative flux for an arbitrary convection in the air can be written as
$j_{e}=\alpha_{e}Sh\frac{Dc_{\rm sat}(1-\mathcal{R}_{H})}{R\rho},$ (1)
where $D$ is the diffusivity of water vapor in air, $c_{\rm sat}$ the mass saturation concentration, $\rho$ is the density of the liquid, and $\alpha_{e}$ a non-dimensional constant. Finally, $Sh$ is the Sherwood number, quantifying the importance of convective evaporation compared to diffusion alone, which will vary depending on the regime of interest. Mass conservation in the cap of the bubble reads:
@@
with $S$ and $P$, the surface and perimeter of the bubble, respectively. Substituting the values of $u$ and $j_{e}$ and using $P/S\sim\ell_{c}/R^{2}$, with $\ell_{c}=\sqrt{\gamma/\rho g}$ the capillary length (Lhuissier and Villermaux,, 2012), we obtain the equation describing the drainage of the cap:
$\frac{\mathrm{d}h}{\mathrm{d}t}+\alpha_{d}\frac{\gamma\ell_{c}}{\mu}\frac{h^{5/2}}{R^{7/2}}+\alpha_{e}Sh\frac{Dc_{\rm sat}\left(1-\mathcal{R}_{H}\right)}{R\rho}=0.$ (3)
We normalize the thickness in the previous equation using the estimated thickness at generation $h_{0}$, and obtain two characteristic timescales $\tau_{d}$ and $\tau_{e}$ representing drainage and evaporation, respectively:
$\tau_{d}=\frac{R^{7/2}\mu}{\gamma\ell_{c}h_{0}^{3/2}}\,{\rm and}\,\tau_{e}=\frac{R\rho h_{0}}{Dc_{\rm sat}},$ (4)
and the drainage equation in dimensionless variables ($\tilde{h}=h/h_{0}$ and $\tilde{t}=t/\tau_{d}$) reads:
$\frac{\mathrm{d}\tilde{h}}{\mathrm{d}\tilde{t}}+\alpha_{d}\tilde{h}^{5/2}+\alpha_{e}\frac{\tau_{d}}{\tau_{e}}Sh\left(1-\mathcal{R}_{H}\right)=0.$ (5)
@@
This model is similar to the one employed by Poulain and Bourouiba, (2018) with the difference that we explicitly specify the evaporative rate in several conditions. The value of the Sherwood number depends on the Reynolds number in the air and the relative humidity. It can be divided into natural and forced convection depending on the value of the Reynolds number. At low Reynolds numbers (i.e. the two lower wind cases of our study), the evaporation is dominated by natural convection. In that case, the evaporation of the surface of the bath creates a plume of lighter high humidity air that rises in the center of the bath. The relevant dimensionless parameter in this problem is the Grashof number comparing evaporation-induced buoyancy and viscous dissipation: $Gr=\Delta\rho_{a}/\rho_{a}\times g\mathcal{L}^{3}/\nu_{a}^{2}$, with $\rho_{a}(\mathcal{R}_{H})$ the density of the air, $\mathcal{L}$ a typical length scale of the bath surface and $\nu_{a}$ the kinematic viscosity of air. The Grashof number can be simplified into $Gr\approx(1-\mathcal{R}_{H})Gr_{0}$ with $Gr_{0}$ the maximal value of the Grashof number, occurring with a perfectly dry air in the chamber (Aurégan and Deike,, 2025). In our setup the Grashof number is typically $10^{4}$. (Dollet and Boulogne,, 2017) have measured a scaling for the Sherwood number associated with the evaporation of the flat surface of the bath in this situation:
$Sh_{N}=\frac{\pi}{2}\left(a_{1}Gr^{0.18}+a_{2}\right),$ (6)
with $a_{1}=0.31$ and $a_{2}=0.48$ two dimensionless constants. Note that the theory in the same reference predicts a scaling with the Grashof number with an exponent $1/5$. In order to use the experimentally fitted values for $a_{1}$ and $a_{2}$, we also use the fitted exponent of $0.18$ (fitted in the case of a circular bath while ours is square). It should be noted that the theory gives a thickness of the boundary layer that is developing above the surface scaling as $\delta_{\mathcal{R}_{H}}\sim R/Gr^{1/5}\approx 5$ mm. The bubbles are therefore fully inside of the concentration boundary layer and expected to evaporate with a similar Sherwood number as the surface of the bath.
@@
At higher Reynolds numbers, the airflow above the bath is turbulent and dominates the evaporation. The rate at which bubbles evaporate is limited by the rate of the renewal of dry air elements close to the surface, which depends on the Reynolds number. The exact form of the relationship between the Sherwood and the Reynolds number depends on the properties of the flow, the geometry of the problem and can also depend on the interfacial properties (Theofanous et al.,, 1976; Turney and Banerjee,, 2013; Wissink et al.,, 2017). This relation can be found by using Higbie penetration theory (Higbie,, 1935; Farsoiya et al.,, 2023): the mass transfer velocity $k_{e}$ is given by $k_{e}\propto\sqrt{D/\tau}$, with $\tau$ the timescale at which air surface elements are renewed. If the relevant timescale for surface renewal is the integral timescale $\tau_{\rm int}=L_{\rm int}/\left|\mathbf{u}^{\prime}\right|$, then the Sherwood number $Sh_{F}=k_{e}L_{\rm int}/D$ reads:
$Sh_{F}\propto Sc^{1/2}Re^{1/2}.$ (7)
@@
If however, the relevant timescale is given by the smallest scales of the turbulent motion, then $\tau\sim(\nu/\varepsilon)^{1/4}$, and we get:
$Sh_{F}\propto Sc^{1/2}Re^{3/4}.$ (8)
@@
In the following, we use a Sherwood number combining the effects of natural and forced evaporation. The forced convection scaling based on the integral timescale is expected to be valid at low Reynolds numbers, while the one based on the small turbulent scales is expected to be valid at higher Reynolds numbers (Theofanous et al.,, 1976). In the following, we will therefore assume that the relevant timescale is the integral timescale. Finally, the Sherwood number can be written as the sum of a term dependent on $1-\mathcal{R}_{H}$, a constant term and term dependent on the Reynolds number:
$Sh=Sh_{N}+Sh_{F}=\frac{\pi}{2}\left(a_{1}Gr^{0.18}+a_{2}\right)+Sc^{1/2}Re^{1/2}.$ (9)
At early times, the evaporative term in Eq. (5) is always negligible compared to the capillary pressure drainage term ($\tau_{d}/\tau_{e}\approx 10^{-4}$, while the dimensionless thickness is of order one), allowing us to recover the classical drainage law for surface bubbles $\tilde{h}\propto\tilde{t}^{-2/3}$. Evaporation becomes the main draining mechanism when the third term becomes of the same order of magnitude as the second one i.e. $\tilde{h}^{5/2}\sim\tau_{d}/\tau_{e}Sh(1-\mathcal{R}_{h})$. Using the early time scaling of the thickness over time $\tilde{h}\propto\tilde{t}^{-2/3}$, this gives the typical dimensionless time $t_{c}/\tau_{d}$ when evaporation becomes the main draining mechanism:
$t_{c}/\tau_{d}=\left(\frac{\tau_{d}}{\tau_{e}}Sh(1-\mathcal{R}_{H})\right)^{-3/5}.$ (10)
@@
Figure 5: Thickness as function of dimensionless time. Markers indicate relative humidity and color shows the dimensionless evaporative rate. Full lines are solutions of Eq. (5), with $\alpha_{d}=0.5$ and $\alpha_{e}=30$. The data shown was obtained with a solution of 200 µmol/L of SDS.
@@
We validate this model only taking into account wind speed through evaporation using experimental data. We analyze the effect of humidity and wind on bubble drainage by measuring the thickness of the cap over time. To do so, we use a high speed camera to record the bursting event. This allows us to track the retraction of the rim as a hole opens in the cap. The thickness is then accessible using the classical Taylor-Culick retraction velocity: $V_{\textsc{TC}}=\sqrt{2\gamma/(\rho h)}$ (Lhuissier and Villermaux,, 2012; Poulain et al.,, 2018; Shaw and Deike,, 2024). We repeat this process for several humidity values and for about 20 bubbles each time. We estimated the thickness at generation by using an average over the earliest bursting events we recorded (with a lifetime shorter than 1 second), and found this initial thickness to be independent of wind speed and humidity and equal to $h_{0}=12$ µm.
@@
The result is shown in Fig 5 using an SDS solution (200 µmol/L): the thickness is normalized by the initial thickness $h_{0}$ and the time is normalized using $\tau_{d}$. For $t/\tau_{d}\lesssim 100$, we find that all the data for different evaporative rates (humidity or wind speed) collapse on the same trend, close to $h\propto t^{-2/3}$. Finding that this classical result is unchanged shows that the strong mixing of the cap at high wind speed actually has little effect on the drainage of the bubbles. Further, the motion of the marginal regeneration patches (their rise velocity or the size at which they detach from the foot of the bubble) does not seem to have a noticeable effect on the drainage in this case.
@@
After a time which is typically of order $t_{c}$ (or $t/\tau_{d}\approx 100$), we see the cases at different evaporative rates separate and the drainage becomes faster. The thickness now decreases linearly with time, with the cases with strong evaporation (yellow) decreasing faster and earlier than the cases with little to no evaporation (black). We can quantitatively predict this behavior by integration of Eq. (5) over time combined with the Sherwood number given by Eq. (9) by fitting the numerical prefactors $\alpha_{d}$ and $\alpha_{e}$ on the data. We see a good agreement with our measurements, further validating the fact that the main draining mechanisms are i) viscous limited flow through the foot of the bubble and ii) direct evaporation from the cap. The numerical value of the prefactors we obtained are $\alpha_{d}=0.5$ and $\alpha_{e}=30$ for the whole dataset. The numerical integration slightly overshoots the cases at high humidity and low wind, possibly because of the hypothesis made in estimating the natural convection Sherwood number.

### v3

@@ 4 Results
Figure 4: (a) Illustrations of thickness variations in the bubble cap using color interferometry. From top to bottom the Reynolds number is 0, 142, and 466. A supplementary movie showing the motion of the patches is available. (b) Lifetime as a function of the integral Reynolds number. Each circle represents a single bubble lifetime, the boxplot adds additional statistical information: (green line) median, (orange dashed line) mean, (box width) interquartile range, (whiskers) furthest data point within 1.5 times the interquartile range from the box. The data corresponding to a fan RPM of 50 or $Re=16$ has been slightly shifted for clarity. The data presented is obtained with 200 µmol/L of SDS and at a relative humidity of 50 %.
@@ 4.1 Qualitative observations
The presence of wind strongly disrupts the classical picture of the motion of the patches in the bubble cap (Illustrated with three panels of Fig. 4 (a), see also the supplementary movie). In these frames white light creates interference patterns within the bubble cap and provides a view of thickness variations. In the top panel (no wind) we see patches generated at the foot of the bubble through marginal regeneration and rising in the cap. These patches allow the exchange of fluid between the bubble cap and the bath and are responsible for the drainage of the cap. They are generated through an instability of the pinch at the foot of the bubble (Lhuissier and Villermaux,, 2012; Gros et al.,, 2021; Trégouët and Cantat,, 2021). These patches are thinner than the rest of the film and therefore buoyant, and they detach from the pinch at the foot when they are large enough. Each patch generated therefore corresponds to a loss of volume in the bubble cap, and it has been shown that the volume change due to generation of these patches exactly matches the viscous limited flow out of the cap (Miguet et al.,, 2021). The motion of these patches in the cap can also be taken into account to model flows induced by surface tension gradients either due to evaporation or heating for instance (Poulain et al.,, 2018).
This picture is largely modified by the presence of wind (middle and bottom panel, and supplementary movie). Instead of rising straight up because of buoyancy, the patches are entrained by the wind in the air close to the cap. They are therefore constantly being deformed and pushed around by the strongly intermittent wind close to the bubble. In the middle panel we can still discern the patches, but some of them have a mostly horizontal velocity instead of vertical and their shape is quite different from the classical picture. In the bottom panel (largest wind forcing) the cap is almost completely mixed with the patches now taking the form of horizontal bands in the cap. In the case with the lowest wind forcing ($Re=16$, not shown), the motion of the patches varies between being very similar to the no wind case and being entrained by the wind as an eddy reaches the proximity of the bath surface. We have verified that the presence or absence of the overflow has no significant effect on the bubble cap images or the conclusions above.
@@
Wind strongly reduces the mean lifetime of surface bubbles in solutions of pure surfactants. For a given solution and relative humidity, the lifetime can vary by a factor up to 5 only due to the intensity of the wind forcing. Fig. 4 (b) illustrates this effect, (with 200 µmol/L of SDS and a relative humidity of 50%) where we can see several effects of the presence of wind: each point is an individual bubble lifetime measurement and the box plot shows the most important metrics of the distribution. The mean lifetime (orange horizontal lines) monotonously decreases as the Reynolds number increases, and so does the spread around the mean. The reduction in the mean lifetime may a priori be due to many effects: early bursting due to a perturbation from the wind, increased drainage through direct evaporation from the cap or increased flow through the bubble foot because of the strong cap mixing for instance. We note here that Burger and Blanchard, (1983) found an opposite trend of increasing lifetime when increasing wind over the surface with a limited dataset. The reason for this opposite trend is unknown. In the following we derive a drainage model taking into account only direct evaporation and show that it is sufficient to explain most of the observed variation of lifetime.
@@ 4.2 Drainage
We model the effect of the wind on bubble lifetime by taking into account its effect on the evaporative rate. The effect of the wind-induced motion in the cap is not considered in the following. There are two mechanisms responsible for the drainage of the cap of the bubbles: viscosity limited flow from the cap towards the bath due to the capillary pressure through the foot of the bubble and direct evaporation from the cap. Following the derivation from previous studies (Lhuissier and Villermaux,, 2012; Poulain et al.,, 2018; Miguet et al., 2020b, ), the velocity through the foot of the bubble reads $u=\alpha_{d}(\gamma/\mu)(h/R)^{3/2}$, with $\gamma$ and $\mu$ the surface tension and viscosity of the liquid, respectively. $\alpha_{d}$ is a non-dimensional constant of order unity. The evaporative flux $j_{e}$ for an arbitrary convection in the air can be written as $j_{e}=\mathrm{Sh}j_{\rm diff}$, with $j_{\rm diff}$ the flux due to diffusion alone, and $\mathrm{Sh}$ the Sherwood number quantifying the importance of convective evaporation compared to diffusion alone. Replacing $j_{\rm diff}$ with its expression gives:
$j_{e}=\mathrm{Sh}\frac{Dc_{\rm sat}(1-\mathcal{R}_{H})}{R\rho},$ (1)
where $D$ is the diffusivity of water vapor in air, $c_{\rm sat}$ the mass saturation concentration, $\rho$ is the density of the liquid. Mass conservation in the cap of the bubble reads:
@@
where $h$ is the thickness of the cap, and $S$ and $P$ are the surface and perimeter of the bubble, respectively. Substituting the values of $u$ and $j_{e}$ and using $P/S\sim\ell_{c}/R^{2}$, with $\ell_{c}=\sqrt{\gamma/\rho g}$ the capillary length (Lhuissier and Villermaux,, 2012), we obtain the equation describing the drainage of the cap:
$\frac{\mathrm{d}h}{\mathrm{d}t}+\alpha_{d}\frac{\gamma\ell_{c}}{\mu}\frac{h^{5/2}}{R^{7/2}}+\mathrm{Sh}\frac{Dc_{\rm sat}\left(1-\mathcal{R}_{H}\right)}{R\rho}=0.$ (3)
@@
The value of the Sherwood number depends on the Reynolds number in the air and the relative humidity. It can be divided into natural and forced convection depending on the value of the Reynolds number. At low Reynolds numbers (i.e. the two lower wind cases of our study), the evaporation is dominated by natural convection. In that case, the evaporation of the surface of the bath creates a plume of lighter high humidity air that rises in the center of the bath. The relevant dimensionless parameter in this problem is the Grashof number comparing evaporation-induced buoyancy and viscous dissipation:
$Gr=\left|\frac{\rho_{a}(\mathcal{R}_{H}=100\%)-\rho_{a}(\mathcal{R}_{H})}{\rho_{a}(\mathcal{R}_{H})}\right|\frac{g\mathcal{L}^{3}}{\nu_{a}^{2}},$ (4)
with $\rho_{a}(\mathcal{R}_{H})$ the density of the air, $\mathcal{L}$ a typical length scale of the bath surface and $\nu_{a}$ the kinematic viscosity of air. The Grashof number can be simplified into $Gr\approx(1-\mathcal{R}_{H})Gr_{0}$ with $Gr_{0}$ the maximal value of the Grashof number, occurring with a perfectly dry air in the chamber (Aurégan and Deike,, 2025). In our setup the Grashof number is typically $10^{4}$. (Dollet and Boulogne,, 2017) have derived an expression for the flux over the surface of the bath: $j_{e}=Dc_{\rm sat}(1-\mathcal{R}_{H})Gr^{1/5}/\mathcal{L}$. Assuming that the flux from the bubble cap is the same one as from the bath’s surface, we obtain the Sherwood number associated with natural convection:
$\mathrm{Sh}_{N}\propto\frac{R}{\mathcal{L}}Gr^{1/5}.$ (5)
Note that in deriving the above equation, we have neglected the effects from the edge of the bath or the central plume similarly to Miguet et al., 2020b .
@@
At higher Reynolds numbers, the airflow above the bath is turbulent and dominates the evaporation. The rate at which bubbles evaporate is limited by the rate of the renewal of dry air elements close to the surface, which depends on the Reynolds number. The exact form of the relationship between the Sherwood and the Reynolds number depends on the properties of the flow, the geometry of the problem and can also depend on the interfacial properties (Theofanous et al.,, 1976; Turney and Banerjee,, 2013; Wissink et al.,, 2017). This relation can be found by using Higbie penetration theory (Higbie,, 1935; Farsoiya et al.,, 2023): the mass transfer velocity $k_{e}$ is given by $k_{e}\propto\sqrt{D/\tau}$, with $\tau$ the timescale at which air surface elements are renewed. The mass transfer velocity is then linked to the evaporation flux in our case through $j_{e}=k_{e}c_{\rm sat}(1-\mathcal{R}_{H})/\rho$, and therefore the Sherwood number associated with forced convection is $\mathrm{Sh}_{F}=k_{e}R/D$. If the relevant timescale for surface renewal is the integral timescale $\tau_{\rm int}=L_{\rm int}/\left|\mathbf{u}^{\prime}\right|$, then the Sherwood number reads:
$\mathrm{Sh}_{F}\propto\frac{R}{L_{\rm int}}Sc^{1/2}Re^{1/2},$ (6)
@@
with $Sc=\nu_{a}/D$ the Schmidt number. If however, the relevant timescale is given by the smallest scales of the turbulent motion, then $\tau\sim(\nu_{a}/\varepsilon)^{1/4}$, and we get:
$\mathrm{Sh}_{F}\propto\frac{R}{L_{\rm int}}Sc^{1/2}Re^{3/4}.$ (7)
@@
In the following, we use a Sherwood number combining the effects of natural and forced evaporation. The forced convection scaling based on the integral timescale is expected to be valid at low Reynolds numbers, while the one based on the small turbulent scales is expected to be valid at higher Reynolds numbers (Theofanous et al.,, 1976). We will therefore assume that the relevant timescale is the integral timescale.
Finally, to continuously cross from a purely convective regime at $Re=0$ to a regime dominated by forced convection at high $Re$, we make the hypothesis that the total Sherwood number is the of sum the two contributions, with an added numerical prefactor $\alpha_{e}$ of order unity as the expressions above are only scaling laws. As a consequence, the Sherwood number can be written as the sum of a term dependent on $1-\mathcal{R}_{H}$, and term dependent on the Reynolds number:
$\mathrm{Sh}=\alpha_{e}\left(\mathrm{Sh}_{N}+\mathrm{Sh}_{F}\right)=\alpha_{e}\left(\frac{R}{\mathcal{L}}Gr^{1/5}+\frac{R}{L_{\rm int}}Sc^{1/2}Re^{1/2}\right).$ (8)
Typical values are 0.3 to 0.5 for $\mathrm{Sh}_{N}$ and 0.5 to 2 for $\mathrm{Sh}_{F}$, we also use the average value $L_{\rm int}=2$ cm in the following.
We obtain the characteristic length and time scales of the problem by noticing that at early times drainage through the foot of the bubble dominates, allowing us to recover the classical drainage law for surface bubbles: $h\propto(\gamma\ell_{c}t/(\mu R^{7/2}))^{-2/3}$ (Lhuissier and Villermaux,, 2012). Evaporation becomes the main draining mechanism when the two draining terms are of the same order of magnitude: $\alpha_{d}\gamma\ell_{c}h^{5/2}/(\mu R^{7/2})=\mathrm{Sh}\left(1-\mathcal{R}_{H}\right)Dc_{\rm sat}/(R\rho)$. This equation allows us to introduce the typical thickness where evaporation becomes the main draining mechanism $h^{\star}$, and using the early time drainage law, the associated typical timescale $t_{c}$:
$h^{\star}=R\left(\frac{\mu Dc_{\rm sat}}{\rho\gamma\ell_{c}}\right)^{2/5}\,{\rm and}\,t_{c}=\left(\frac{R^{2}\rho}{Dc_{\rm sat}}\right)^{3/5}\left(\frac{\mu R^{2}}{\gamma\ell_{c}}\right)^{2/5}$ (9)
We finally obtain the dimensionless drainage equation by normalizing Eq. (3) using $\tilde{h}=h/h^{\star}$ and $\tilde{t}=t/t_{c}$:
$\frac{\mathrm{d}\tilde{h}}{\mathrm{d}\tilde{t}}+\alpha_{d}\tilde{h}^{5/2}+\alpha_{e}\tilde{j_{e}}=0,$ (10)
with $\tilde{j_{e}}=\left(\mathrm{Sh}_{N}+\mathrm{Sh}_{F}\right)\left(1-\mathcal{R}_{H}\right)$ the dimensionless evaporative flux. This model is similar to the one employed by Poulain and Bourouiba, (2018) with the difference that we explicitly estimate the dependence of evaporative rate with wind speed and humidity.
@@
Figure 5: Cap thickness as a function of time in dimensionless units. Markers indicate relative humidity and color shows the dimensionless evaporative rate $\tilde{j_{e}}=\left(\mathrm{Sh}_{N}+\mathrm{Sh}_{F}\right)\left(1-\mathcal{R}_{H}\right)$. Full lines are solutions of Eq. (10), with $\alpha_{d}=1.8$ and $\alpha_{e}=1.2$. The data shown was obtained with a solution of 200 µmol/L of SDS.
@@
We validate this model only taking into account wind speed through evaporation using experimental data. We analyze the effect of humidity and wind on bubble drainage by measuring the thickness of the cap over time. To do so, we record the spontaneous bursting of surface bubbles using a high speed camera. This allows us to track the retraction of the rim as a hole opens in the cap. The thickness is then accessible using the classical Taylor-Culick retraction velocity: $V_{\textsc{TC}}=\sqrt{2\gamma/(\rho h)}$ (Lhuissier and Villermaux,, 2012; Poulain et al.,, 2018; Shaw and Deike,, 2024). We repeat this process for several humidity values and for about 20 bubbles each time.
@@
The result is shown in Fig 5 using an SDS solution (200 µmol/L): the thickness is normalized by the units defined in Eq. (9). Different markers correspond to different relative humidities and color indicates the dimensionless evaporative rate $\tilde{j_{e}}=\mathrm{Sh}\left(1-\mathcal{R}_{H}\right)$. For $t/t_{c}\lesssim 0.4$, we find that all the data for different evaporative rates (humidity or wind speed) indeed collapse on the same trend, close to $h\propto t^{-2/3}$. Finding that this classical result is unchanged shows that the strong mixing of the cap at high wind speed actually has little effect on the drainage of the bubbles. Further, the motion of the marginal regeneration patches (their rise velocity or the size at which they detach from the foot of the bubble) does not seem to have a noticeable effect on the drainage in this case.
After a time which is typically of order $0.4\times t_{c}$, we see the cases at different evaporative rates separate and the drainage becomes faster. The thickness now decreases linearly with time, with the cases with strong evaporation (yellow) decreasing faster and earlier than the cases with little to no evaporation (black). As evaporation becomes dominant, the location where the bursting event is initiated goes from being homogeneously distributed in the cap to being more likely close to the top (see App C).
@@
We can quantitatively predict the drainage behavior by integrating Eq. (10) over time combined with the Sherwood number given by Eq. (8) by fitting the numerical prefactors $\alpha_{d}$ and $\alpha_{e}$ on the data. Note that in order to do so, we need to estimate the thickness at generation $h_{0}$. We estimated this value by using an average over the earliest bursting events we recorded (with a lifetime shorter than 1 second), and found this initial thickness to be independent of wind speed and humidity and equal to $h_{0}=12$ µm. All other physical parameters of the model are either measured or computed from the properties of the solution. We see a good agreement with our measurements, further validating the fact that the main draining mechanisms are i) viscous limited flow through the foot of the bubble and ii) direct evaporation from the cap. The numerical value of the prefactors we obtained are $\alpha_{d}=1.8$ and $\alpha_{e}=1.2$ for the whole dataset. The numerical integration slightly overshoots the cases at high humidity and low wind, possibly because of the hypothesis made in estimating the natural convection Sherwood number.